OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Numerical Answers Expected!

OpenStudy (anonymous):

plug in y = 125,000 and solve for x

OpenStudy (anonymous):

x-62500 = 125000*0.25 x-62500 = 31250 x = 31250+62500 x = 93750

OpenStudy (anonymous):

thank you guys soooo much :D))